L. 12) Here are three oscillator differential equations: y[t]0.1 y [t] + 5.8 y[t] = 0 » y[t] +5.8 y[t] = 0 , y t]14 y t]5.8Here are three oscillator differential equations: y''[t] + 0.1 y'[t] + 5.8 y[t] = 0 , y''[t] + 5.8 y[t] = 0 , y''[t] + 1.4 y'[t] + 5.8 y[t] = 0 . One is a differential equation of an undamped oscillator; the other two are differential equations of damped oscillators. Which is which? Of the two damped oscillator differential equations, which one is the more heavily damped?

PLEASE INCLUDE EXPLANATION!!!
Answer #1

Given three equations yCt) + 0.1 y( t) + 5-8 yC4) = 0 y (t) + 5.87%) 0 yCt) + 1.4 4(4) +5.8 4(4):0. Convert them to lapsc + S 8 9 1) = 0 3 4 5 +5.87(5)=0 (5 + 5.8) 4 (s)=0 For this y=0 Undamped This is an case. B y(t) + 1.4 yrt) + 5.8 g (t)